A Complete Bibliography of Publications in The European Physical Journal H: Historical Perspectives on Contemporary Physics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

01 October 2024 Version 1.29

Title word cross-reference

3 [FV14]. ^{+} [BP11]. ^{−} [BP11]. E = mc^{2} [Bou13]. e^{+}e^{−} [BW22]. H [Bad11, Shi24, vN10]. Λ [Nov18]. p [DJ13]. q [DJ13]. Υ [SM11]. W [Tan16].

-Curve [Shi24]. -theorem [Bad11, vN10].

1911 [Str11]. 1920s [CDM22, PP16]. 1930s [DEN24, GS24]. 1950s [Old23]. 1953 [Cro11]. 1957 [Fre10]. 1960s [CDM22]. 1966 [Ell14]. 19th [BLR20, Sch14b].

20th [BLR20, DFP18].

50th [AV24]. 51 [CB15].

60 [Kik18]. 60-year-old [GK23]. 60s [Sch10].

70s [Sch10]. 76 [FBB14].

81 [PMG23].

abandoned [OMNM14]. abstract [Bor13]. abundances [Tri10]. accelerator [DLL19]. accelerators [Sch12]. achieve [Kik18]. achievements [BW22, Haa10b]. across [Nar18]. action [Ish17, PB17]. Adams [Ell14]. Addressing [BCC^{+}24]. after [Des18, LP24, ZLK11]. against [Hel21]. age [Tip13]. air [Dah15, KW12]. Alder [BC18]. algebra [Pie11]. algebras [Fre10, Haa10b]. allgemeinen [Ein31]. almost [FV14]. Alpher [AM23]. American [AV24, RBM14]. amplifier [dDD^{+}19]. Analogue [AJ23]. Analysis [PB17, BH17, OM14, OM22, PAP21]. Anatoly [GK23]. angle [DLL16]. Anglo [RBM14]. anniversary [AV24]. Annotation [FBB14]. anode [Dit16]. anti [Dit15b]. anti-Dirac [Dit15b]. antiproton [Rub12]. antiquity [Lev16]. apple [Gor24b]. application [CM12]. applied [RMS21]. approach [BR13]. approaches [Sal22]. approximation [AV24, Ste16]. Arguments [Hel21]. Arrhenius [Kra13a]. Aspects [Gal13]. assumption [Fic13]. assumptions [Kre15]. astrometry [Per12]. Astronomer [RBM14]. astronomy [LW12, Spi12b]. astrophysical [NHK^{+}22]. astrophysics [Bon17, KKS12, MC18]. asymptotic [Haa10a]. atom [Eck14]. Atomic [Cro11, Bai13, GS22, Kra11]. atoms [Hol21]. attempt [Sch19, Sch20]. Attempts [JN14, La 22]. axiom [Haa10a].

b [CB15]. back [Haa10b]. background [MM24]. Bagnères [Cro11]. Balmer [Som14b]. Bang [Pee14]. Bartoli [BLR20]. based [Dua12]. basis [vS21]. be [Sch19, Sch20]. beam [CM12]. beams [DLL19, Per16]. became [Sch19, Sch20]. before [Bon17, Tri10]. beginning [MC18]. beginnings [DB14]. behavior [GLTZ10]. belated [HR23]. Berni [BC18]. Bethe [FBB14]. between [AV24, BLR20, Dar23, DD13, DEN24, Hol21, PMG23, Rad13, Sch19, Sch20]. Beyond [dA22, LJ15]. Beziehungen [PMG23]. Big [Pee14]. Bigorre [Cro11]. binary [Ken17]. Binder [Mar19]. biochemistry [MRC22]. Birkeland [Kra13a]. Birth [Cro11, Piz16, DEN24, Eck10, Fre10, LP24]. blackboards [VT22]. blossoming [LR20]. bodies [SKL14]. Bohm [Old23]. Bohr [Eck14, Gor24a, Kra11, Lau24, PP16]. Boltzmann [Bad11, Shi24]. Bonn [Pau13]. Born [AV24, HR23]. boson [Bor15]. Bothe [Fic13]. boundary [Hol21]. bounds [MP13]. Braking [Bet14, FBB14]. branches [Kre11]. breakup [MM24]. Bremsformel [Bet32]. bridge [Sch19, Sch20]. bridge-building [Sch19, Sch20]. brief [CDM22, Des18, JN14, MM24]. Bronstein [Gor24a]. Brownian [Gen20]. Bruno [BP11]. Bryce [BH17]. bubble [Per16]. bubbles [MM24]. building [Sch19, Sch20]. ca [Kra16]. campaign [Sch14b]. canonical [RB15]. career [TS19]. Carlo [GS24]. Cartan [Sch20, Sch19]. case [BLR20, CD11, PI22]. Cauchy [FV14]. caused [HS16]. CCBA [AV24]. centenary [dA22]. centennial [OONM17]. centre [RMD23]. centuries [BLR20]. Century [Nov18, Bor13, DFP18, Haa10c, Sch14b]. CERN [Hüb12, Per16, Pla12, Ric12, Sch12]. certain [PMG23]. chamber [Per16]. Chandra [RBM14]. channels [AV24]. chaos [Esc18]. charged [SKL14]. Charles [Joh20]. Charm [MB17a, Dor12]. chemical [Tri10]. Clarifying [DD13]. Classical [SKL14, GG19, Hel21, Kre15, LJ15, PP19, PMG23, vS21]. close [DEN24]. cofounder [Dit15a]. collaboration [AM23]. collaborative [SGD20]. collapse [Alm20]. College [Rob19]. collider [BW22, BP11, Tan18]. colliders [Ric12, Rub12]. commented [Shi24]. commitment [Haa10c]. Como [Mar19]. compact [Bon17]. composite [Haa10a]. Computational [Mar18, HM24, LH19, MRC22, SGD20]. computer [AV24]. computers [Hol21]. concept [DEN24, Fre10, Gen20, GS22]. Conceptual [Kra11]. concrete [Smi23, Smi24]. condensed [DB14, Kar22]. conditions [Kre15]. conducted [AV24]. conference [LP24, Cro11]. confinement [HM18, Wag18]. Congress [Str11]. Connection [RBM14]. constant [Fra13, JN14, OONM18]. constrained [SS17]. Constructive [Shl12]. contemporary [FGV17]. context [LR20]. continued [Pla12]. continuum [Hol21]. Contour [DPP^{+}12]. contributed [Tan18]. contributions [Sch10]. controversy [GK23, Ken17]. conundrum [Hic12, Sch13]. conversion [Nus14]. cooling [CM12]. coordinates [Nie22]. Copenhagen [BH17]. Corbino [BLR20]. cornerstone [Kre15]. Correction [CF24b, HMS23a, OM22, Smi24, IN16]. correspondences [PP16]. Cosmic [Cro11, Spi12a, CD11, KW12, Mül12, OM14, OM22, WW12]. cosmical [Kra13a]. cosmogonies [Kra13a]. Cosmological [Ein31, AM23, OONM18]. cosmology [Hel21, Nar18, OONM15, VT22]. cosmos [OMNM14, OONM15]. Cosserats [Sch20, Sch19]. could [Gor24b]. coupled [AV24]. CP [Cro12]. CPT [BdV22]. CRC [AV24]. credible [Ste16]. criteria [Shl12]. critical [Ble12, SM11]. criticality [Mar19]. crystals [SCS22]. cultural [Ben10]. Curie [Rad13]. current [Pie11]. curvature [Sch19, Sch20]. Curve [Shi24].

D [FV14]. dark [OONM18]. Darwin [Joh20]. dawn [Fic13]. day [Lev16]. days [Pic12]. death [dA22]. decades [Nar18]. decelerated [HS16]. dedicated [TWZ19]. Definitions [MPP^{+}12]. deflection [Sch14a]. Department [Dor13]. detectors [Piz16]. determination [JN14]. determine [Sch14b]. determinism [vS21]. deterministic [vS21]. detonation [Kre11]. develop [TWZ19]. development [DEN24, HM24, Ina15, NHK^{+}22, Pla12, Tan18, Wei21, ZLK11]. developments [Piz16]. DeWitt [BH17]. diagrams [Dar19, HM24]. Dicke [Pee17]. different [CM12]. Dirac [Dit15b, PI22]. Direct [Mül12]. discipline [Goe17]. discover [Gor24b]. discovered [Smi13]. Discovery [Pee14, AK11, CD11, CB15, Cro12, Eck12, Pen13, Per14, Söd10, Ste16, SM11].

Discussion [Haa10a, Rad13]. dispersion [Joh20, LL14, Tal20]. dissident [Old23]. dissipation [Dar23]. do [Kra11]. dozen [Iof13]. drama [Gor24a]. during [AV24]. Dynamical [GS12]. dynamics [FGV17, Kar22, MRC22, SS17]. Dyson [Ble12].

E. [Rad13, Sch19, Sch20]. Early [Bai13, LJ15, Mar18, Per16, WW12, Bad11, Bor15, Haa10b, Kra13b, PP16, Per13, PAP21, Pie11, TS19]. Editorial [Bei10, BG14, BGLR17, BL24, DFP18]. effect [AJ23, Eck12, Smi13]. effective [RV24, Wei21]. effects [Roc23]. Ehrenfest [PP16]. einige [PMG23]. Einstein [BGLR17, Sch20, Nus14, OM14, OMNM14, OONM15, OONM17, OOM21, OM22, Sau19, Sch19, Sch13, VT22]. Elections [Bet32]. electrodynamics [Dar19]. electromagnetic [SKL14]. Electron [Ric12, Pel12]. Electrons [FBB14, Bet14, Kra11]. Elektronen [Bet32]. elements [AM12, Kra13b, PI22, Tri10]. emergence [Kre11, MRC22]. encounters [DEN24]. end [CDM22, HMS23a, HMS23b]. energy [KW12, Kra16, LW12, Mas14, OONM18, Spi12b, TWZ19]. engineering [Kre11]. enhancement [CM13]. ensemble [Ina15]. entanglement [Dua12]. entropies [DD13]. episode [Pen13]. episodes [FP23]. equation [FV14]. equations [Dar10]. Equilibrium [Gal16, Alm20, LJ15]. era [HMS23a, HMS23b, MC18]. ergodic [vN10]. Ergodicity [Gal16]. Erratum [Sch20, Zag13]. erroneous [Ste16]. Essay [Ell14]. estimation [Bis15]. ether [Dar10]. eugenics [Hal24]. Euler [Bis15, FV14]. EUROGRAV [La 22]. Europe [DEN24]. European [Reb18, La 22]. ever [vS21]. Everett [Wu21]. everything [Sin12]. evolution [Nar18, Old23]. examples [Kar22]. exchange [DEN24]. exemplifies [Mou13]. existence [MPP^{+}12]. Exoplanets [MC18]. expanding [Nus14]. expansions [FR24]. experience [CB15]. experiment [AV24, BW22, KKS12, Per13, SBSL^{+}16]. experimental [AM12, Pau13, Pee17]. experiments [Dor13, Dua12, Per16, Piz16, Sch12, Tan18, TWZ19]. exploration [LJ15]. Exploratory [DeW17]. Exploring [Hol21]. extended [Eck14, LJ15, SKL14]. Extensive [KW12].

facets [DM11, Mou13]. falling [Gor24b]. fast [HS16]. Father [RMS21, BP11]. favorite [GS22]. Felici [BLR20]. Fermi [GS22, PI22, Zag11, Zag13]. few [Pic12]. Feynman [Dar19, HM24, Zeh11, dDD^{+}19]. field [ABR23, Dar19, DeW17, Dit15a, DM11, Fic13, FR24, GG19, Haa10a, RV24, Sau19, Sch10, Sch11, Wei21]. field-less [Dar19]. fields [MPP^{+}12, Ros17]. fifth [Fis15]. fight [Kik18]. figure [GS22]. fin [Kra13a]. fin-de-siècle [Kra13a]. fine [JN14, Som14a]. fine-structure [JN14]. first [Des18, Iof13, JN14, LR20, La 22, Sch14a, Ste16, Str11, Tal20]. fission [Kra14, Ste16]. flatness [Hel21]. Florence [LP24]. Florentine [CDM22]. flow [FV14]. fluctuation [Dar23]. fluid [Cra12, FGV17]. fluids [Bis15, VR17]. force [Fis15, SKL14]. forces [DEN24]. forgotten [FH22, FV14, OM14, OM22]. formative [Kon20]. Formula [FBB14, Bet14, Ken17]. formulation [FV14]. formulations [SKL14]. Förster [Mas14]. foundational [Sch13]. foundations [DJ13, FR24]. fountain [Bis15]. four [Mou13]. Fractals [IN16]. France [LH19]. free [Kra11, Pel12]. free-electron [Pel12]. French [LPA18]. frequency [Lev16]. FRET [Mas14]. friction [Bis15]. Frisch [Ree24]. Fritz [Bou13]. frontline [BT21]. full [Kar14]. function [Wu21]. funding [TWZ19]. Fürth [PMG23]. fusion [Esc18, Kik18].

G [RBM14]. galactic [Mül12]. Galton [Joh20]. gamma [LW12, Spi12a]. gamma-ray [LW12]. Gamow [AM23]. gas [MM24]. General [Goe17, Rob19, SS17, Sal22, VR17, Ein31]. genesis [Blu14, BdV22, RMD23, RB15]. geography [Des18]. geometry [Ben10, Haw14]. Geon [Ric18]. George [AM23]. Gerlach [SBSL^{+}16]. German [Bet32, Ein31]. Germany [Goe17]. Geschwindigkeit [Bet32]. GeV [SM11]. Gian [GS24]. Gian-Carlo [GS24]. Gibbs [MPP^{+}12]. gluon [AK11, Söd10, SM11]. Gravitation [BGLR17, Rob19, Goe17, RV24]. Gravitational [La 22, Alm20, DeW17, Piz16, Roc23]. Gravity [PR18, AJ23, BH17, Gor24a, Gor24b, Pee17]. group [Ble12]. growth [Goe17]. Gullstrand [Nie22].

Hacking [Sch15]. hadron [Tan18]. Hahn [Ste16]. half [Haa10c]. Hamilton [Sal22]. Hamiltonian [Esc18, SS17]. Hankel [FGV17, VR17]. Hans [FBB14, SCS22]. happened [Pee14]. hard [Tan18]. hard-scattering [Tan18]. Hasenöhrl [Bou13]. Hawking [AJ23, Ell14]. Heidelberg [Wei15]. heights [Bis15]. Heinrich [dA22]. helped [Gor24b]. heritage [GK23]. Hermann [FGV17, VR17]. heuristic [Fic13]. hierarchical [Ble12, Ble12]. Higgs [Bor15, SZ12]. high [KW12, LW12, Spi12b, TWZ19]. high-energy [KW12, Spi12b, TWZ19]. Historical [OOM21, AK11, AJ23, Dor12, Gal16, JN14, KW12, LPA18, MM24, vS21]. History [CM12, DLL19, Eck12, RD11, Bad11, BCT21, BW22, CDM22, Dar23, Des18, Fis15, FP23, Gen20, Gor24a, GS22, Hol21, IN16, Iof13, Kik18, Lau24, LJ15, Les24, Lev16, MAE^{+}19, Pel12, Per14, Per12, Pie11, Sal22, Wag18, WW12, Bil13]. honor [DS10]. hot [LPA18, Pee14]. Hughes [Eck12]. Hugoniot [Kre15]. hundred [BT21, OONM18]. hydrodynamic [Eck10]. Hydrogen [Som14a]. Hydrogen-like [Som14a]. hypothesis [Kra16, Tri10].

ideal [Kre15]. ideas [Gor24a]. II [LP24]. illuminated [Lor19]. image [vS21]. impact [AK11]. implanted [Kar14]. important [Kre15]. improved [Wag18]. incompressible [FV14]. Inconsistency [Zag11, Zag13]. Indian [RMS21]. Indistinguishable [PI22]. Inertia [Lan14, Pfi14]. infinitely [GS12]. infrared [dA22]. initial [Piz16]. insomnia [Bad11]. institutional [Ben10]. integration [Sch12]. interactions [Dor12, DLL16, LPA18, Per16]. interference [CF24b, CF24a]. Interferometric [La 22]. Internal [Kar22]. international [LP24, Sch12, Sch14b]. internationalism [CD11]. internment [Hal24]. interplay [AV24]. interpretation [RD11, Tal20, Zeh11]. Intersecting [Hüb12]. interview [SD18]. intriguing [dDD^{+}19]. introduction [BGLR17, DFP18]. inverse [BT21]. Ireland [Ben10]. irradiation [HS16]. irreversible [LJ15, MW12]. Ishiwara [PB17]. Ising [FH22]. isotopes [HS16]. ISR [Hüb12]. issue [BGLR17, DFP18]. Italian [BLR20, LR20]. Italy [LR20]. ITEP [Iof13]. ITER [PAP21]. IUCAA [RMD23].

Jack [Tse12]. Jacobi [Sal22]. JADE [BW22]. James [DS10, Dar10, Spe10]. Japan [DEN24]. Japanese [Hal24, Kik18]. JET [Reb18, Bis15]. jets [AK11, AK11]. Joint [Reb18]. Jordan [CF24b, CF24a, Dit15a, DJ13, Sch11, Sch13]. Jorge [Sch10]. journey [ZLK11]. JT [Kik18]. JT-60 [Kik18]. jump [Kre15]. Jun [PB17]. juncture [LP24].

KAGRA [BCC^{+}24]. Kamiokande [KKS12]. Kaplan [NHK^{+}22]. Kelvin [Cra12, Tip13]. King [Rob19]. klassischer [PMG23]. known [Sch19, Sch20]. Korcak [IN16]. Korcak-law [IN16]. Kosmologischen [Ein31]. Kubo [Kon20]. Kurt [Mar19].

LAA [TWZ19]. Laboratories [Eck12]. Laboratory [Iof13]. Ladenburg [Tal20]. Lagrange [Gal13]. Lagrangian [FV14, FGV17]. land [Mar19]. Landau [Gor24a]. Lange [Pfi14]. large [Kik18]. laser [ZLK11]. lasers [Pel12]. late [Bor13]. lattice [MPP^{+}12]. law [CF24b, CF24a, IN16, Lan14, Pfi14]. legacy [Gal13, Sch11]. Léon [SS17]. LEP [Mye12, Pic12, SZ12, Tse12]. less [Dar19]. Lev [Gor24a]. LHC [MB17b, SZ12]. LHCb [BCT21]. life [LB17]. Light [Lor19, RBM14, ABR23, FK19, Roc23]. LIGO [BCC^{+}24]. like [Som14a]. Lille [Fre10]. limit [Kra13b]. lines [Som14a]. link [DD13]. liquid [SCS22, Tan18]. literature [AJ23, BCC^{+}24]. lives [EA24]. Local [Haa10b, FP23, Fre10, Haa10a, LJ15, Sch13]. London [Rob19]. Long [GLTZ10, Mar19, ZLK11]. Long-time [GLTZ10]. longitude [Sch14a]. look [FGV17, Haa10b]. Lord [Cra12]. Lorentz [PP19]. Lorenz [FK19]. lower [MP13]. Ludvig [FK19]. Ludwig [Pfi14]. Luis [GG19]. Lviv [NHK^{+}22].

MacCullagh [Ben10, DS10, Dar10, Spe10]. macroscopic [GLTZ10]. Made [Com29]. magic [Dar19]. magnetic [ABR23]. magnetocaloric [Smi13]. magnetrons [Dit16]. many [CM12, GS12]. mass [Bor15]. Masters [BLR20]. materials [Lau24]. mathematical [Haa10c, Ble12, DPP^{+}12, GS12, MPP^{+}12, Shl12, Sin12]. mathematics [LL14, RMS21]. Matrix [HR23, CF24b, CF24a]. matter [DB14, Kar22]. Matvei [Gor24a]. Max [HR23]. Maxwell [Dar10]. meaning [Ish17, PB17]. Measurement [DLL16, BR13, Fra13, Lau24, Sch14a]. measurements [Dua12, FP23, RD11]. mechanical [Bai13]. Mechanics [Mar18, BR13, Cra12, EA24, HR23, LR20, Lau24, LB17, LH19, MRC22, PMG23, vN10, Gal13]. Mechanistic [Bad11]. meeting [BH17]. Memorandum [Ree24]. Memories [TS19, Tse12, Pic12]. mentor [AM23]. metals [Kar14]. meteoric [Kra16]. methods [HM24]. microscopy [CM13]. Millikan [Fra13, Hal24]. Mind [DJ13]. mineral [SCS22]. mis [HR23]. mis-prized [HR23]. missed [Haa10b, Per13]. mission [Kik18]. model [Ble12, Eck14, FH22, OM14, OMNM14, OONM15, OONM17, OOM21, OM22]. modeling [FP23]. models [Bai13, Ble12]. modern [AV24, Kre15, Nar18]. Molecular [SGD20, BC18]. Moore [BT21]. Moscow [Ble12, DPP^{+}12, GS12, MPP^{+}12, Shl12, Sin12]. most [AV24]. motion [Gen20, VR17]. multiplication [CF24b, CF24a]. muon [Kar14]. my [TS19]. mystery [dDD^{+}19].

N. [RMS21]. naissance [Pee17]. name [MUN23]. names [Kra14]. Naples [EA24]. national [LP24]. Nationalism [CD11]. necessary [Blu14]. Neumann [DD13, DJ13, Lau24]. neutral [PAP21]. Neutrino [Bil13, Dor13, Dor12, DLL16, DLL19, KKS12, Per13, Per14, Per16, Spi12b]. neutrinos [Spi12a]. neutron [GS22, GS24]. neutrons [HS16]. Never [DJ13]. Newton [Gor24b]. Niels [Gor24a]. nineteenth [Bor13]. nobelization [HR23]. Nonequilibrium [Gal16]. Nordic [Kra13a]. Note [FP23, PP19]. Nuclear [Wei15, DEN24, Ric18, Ste16]. nucleon [AV24]. numerous [Kre11].

objections [Kra11]. objects [Bon17]. oblateness [RD11]. observables [Fre10, Sal22]. observations [Mül12]. Observatory [La 22]. old [GK23]. One [OONM18]. ongoing [Sch11, Sch13]. ontological [Old23]. open [VT22]. opening [AM12]. operation [Mye12, Pla12]. opportunities [Haa10b]. optical [Dar10, Joh20, Tal20]. organization [HM18]. origin [Bor15, Dua12, KKS12, LH19, Nie22]. origins [PI22, Tri10]. Orthodox [Old23]. oscillations [Per14]. overview [JN14]. Oxford [VT22].

P. [Rad13]. Painlevé [Nie22]. paper [Dit15b, FBB14, PMG23]. paradigm [AM12]. parallax [Sch14b]. particle [Bil13, BP11, Bor15, HMS23a, HMS23b, LPA18, Pau13, PAP21, WW12]. particles [Dor12, GS12, Haa10a]. Pasadena [Tri10]. Pascual [CF24b, CF24a, Dit15a, Sch11]. past [MW12]. Paths [Mas14]. Paul [RB15]. Pauli [Dit15b]. pedestrian [BR13]. Peering [Fri12]. Pegasi [CB15]. Peierls [Ree24]. pendulums [Dah15]. Penrose [BT21]. pentaquark [Hic12]. people [Haa10c]. Percolating [EA24]. perfect [Tan18]. period [AV24]. periodic [Kra13b]. Perry [Tip13]. personal [DB14, Fis15, Hol21, MB17b, Nar18, Pau13, Smi23, Smi24, Wei15]. perspective [AJ23, Gal16, OONM15, Smi23, Smi24]. Perturbative [FR24]. PETRA [BW22]. phenomena [Ble12]. philosophical [OOM21]. photon [Kra14]. physicist [BP11, Kon20, Ric18]. Physics [BLR20, Ble12, Cro11, DPP^{+}12, Dor12, GS12, Iof13, MB17a, MPP^{+}12, Shl12, Sin12, BT21, Ben10, Bor15, CDM22, DB14, DFP18, GS22, GS24, Haa10c, Kra13a, Kra14, Kre11, Kre15, LP24, Les24, Mou13, PP19, Pau13, Pee17, SD18, Sch13, TWZ19, TS19, WW12, Wei15, vS21, Dor13]. Physik [PMG23, FBB14]. pioneering [BC18, FGV17]. Planck [Fra13]. plane [Lor19]. plasma [HM18, Kra14, SD18, DFP18]. plasmas [LPA18]. players [DFP18]. plus [CB15]. PLUTO [SM11]. point [LPA18]. polarisability [MP13]. polarization [Dua12]. portals [VT22]. positive [Kar14]. possible [Blu14]. prehistory [PR18]. presence [ABR23]. present [Lev16, MW12, Sch10]. principles [JN14]. Prize [Ell14]. prized [HR23]. probability [Zag11, Zag13]. Problem [Ein31, BCC^{+}24, BR13, Eck10, Hel21]. problems [Haa10c]. process [Lau24]. processes [MW12]. produced [Dor12]. production [HS16]. project [TWZ19]. projects [CM12, SGD20]. Proof [vN10]. propagation [Lor19]. properties [Haa10a]. protégé [AM23]. protein [MRC22]. Proton [Rub12, Fri12, Pla12]. Proton-antiproton [Rub12]. Prout [Tri10]. PS [Tse12]. pseudopotential [GS22]. Publisher [CF24b]. pulsar [Ken17]. pulsars [Pen13].

QCD [AK11, Smi23, Smi24, Tan18]. QED [Dar19]. QFT [FP23]. quadrupole [Ken17]. Quantenmechanik [PMG23]. quantization [DeW17, RB15, Ros17]. Quantum [PR18, Bai13, BH17, BR13, Dit15a, DM11, Dua12, DJ13, Fic13, FR24, GLTZ10, Gor24a, Ish17, Joh20, LR20, Lau24, PMG23, PB17, PI22, RV24, Sch10, Sch11, Sch13, Sch15, Tal20, Zag11, Zag13, Zeh11, vN10]. quark [AK11, Per13]. quarks [Per16, Smi23, Smi24]. question [Ber20].

R [RMS21]. R. [PMG23, RBM14]. radioactive [Kra16]. radium [HS16]. radius [RD12]. Ralph [AM23]. Raman [Eck12]. random [MPP^{+}12]. Rankine [Kre15]. rapid [ZLK11]. rational [Mou13]. Ray [Cro11, LW12, Pel12]. rays [CD11, KW12, Mül12, Spi12a]. reaction [AV24]. reactions [AV24]. Real [Ber20]. reality [Kre15]. recollection [Dor12, MB17b, Pau13, SM11]. recollections [Wei15]. reconstruction [LP24]. reference [dDD^{+}19]. Reflections [Mou13, OOM21, Old23]. regimes [Wag18]. regions [FP23]. relation [Dar23, Sch13]. relations [LL14, PMG23]. Relativistic [Bet32, FBB14, Bet14, Bon17, DM11]. relativistischer [Bet32]. Relativitätstheorie [Ein31]. relativity [Goe17, PP19, Rob19, Sal22, TS19, Ein31]. relaxation [Kar22]. relevance [Sch10]. reliable [Pla12]. remarkable [Per14]. remembrance [DB14]. Renaissance [BGLR17]. renormalisation [Ble12]. report [BH17]. Research [Eck12, BT21, DeW17, Kik18, NHK^{+}22, RMD23, Sch10, Sch11, Sch13, Spi12a, Wag18]. resistance [Dah15]. Resolution [CM13]. resonance [Mas14]. resonant [Piz16]. response [dDD^{+}19]. results [SM11]. review [AK11, AJ23, AM23, Hel21, KW12, OONM15, OONM17]. revisited [OM14, OM22, SBSL^{+}16]. Revisiting [Ree24]. revival [BW22]. revolution [Sch15]. Rights [RD12]. Rings [Hüb12]. Roald [SD18]. Robert [Hal24, Pee17]. Rome [Dor13, DB14]. roots [LR20]. Rosenfeld [SS17]. route [Dar10]. Routes [Bor13]. Rubens [dA22]. Rudolf [Tal20]. Russia [SD18]. Rutherford [Rad13]. Ryogo [Kon20].

’s [DJ13, BH17]. Sagdeev [SD18]. Samuil [NHK^{+}22]. Santaló [GG19]. Sapienza [Dor13]. scattering [FK19, Tan18, Bet32]. Schottky [FH22]. science [CD11, Kre11]. scientific [BCC^{+}24]. search [Sau19]. Searching [SZ12, dDD^{+}19, Piz16]. seen [Nar18]. self [HM18, SKL14]. self-force [SKL14]. self-organization [HM18]. seminar [Ble12, DPP^{+}12, GS12, MPP^{+}12, Shl12, Sin12]. Sen [RMS21]. series [Som14b]. session [DeW17]. SETI [Pen13]. sheets [Sau19]. Shock [Kre11, Kre15, MM24]. showers [KW12]. siècle [Kra13a]. simulation [BC18, SGD20]. simulations [MRC22]. simultaneity [PP19]. Singularities [Haw14]. Sitter [OONM15, OOM21]. six [Nar18]. slumber [Bad11]. society [AV24]. sogenannte [Shi24]. solar [Kra16, RD11, RD12, Sch14b]. Solvay [Str11]. Some [Haa10c, Haa10b, Kar22]. Sommerfeld [Eck10, Eck14]. source [Kra16]. space [Ben10]. spacetime [FP23, Haw14, Ric18]. special [BGLR17, DFP18, PP19]. speculations [Kra13b, Kra16]. Speed [FBB14, ABR23, Bet14]. sphere [Lor19]. spheres [FK19]. spin [Blu14, Kar22]. split [Dit16]. split-anode [Dit16]. spreading [LR20]. stability [Eck10]. started [Sin12]. State [Ble12, DPP^{+}12, GS12, MPP^{+}12, Shl12, Sin12, OMNM14]. states [Zag11, Zag13]. static [Nus14, OONM17]. Statistical [Ble12, DPP^{+}12, GS12, Mar18, MPP^{+}12, Shl12, Sin12, Bad11, EA24, LB17, LH19, MRC22]. statistics [Blu14, PMG23, PI22]. Statistik [PMG23]. StatPhys [LP24]. steady [OMNM14]. steady-state [OMNM14]. Steinberger [Tse12]. Stellar [Alm20, Bon17]. step [FH22]. Stephen [Ell14]. steps [Per16]. Stern [SBSL^{+}16]. stimulated [Eck12]. stimulus [Kre11]. stochastic [CM12]. Storage [Hüb12]. story [Bor15, Kar14, SD18]. Stracciati [BLR20]. Strassmann [Ste16]. strings [RV24]. Strong [HM18]. structure [Bon17, JN14, Som14a]. students [BLR20]. studied [Kar22]. stunt [OONM18]. Sub [Cro11, Goe17]. sub-discipline [Goe17]. substructure [Per13]. sun [Tip13]. superfluous [OONM18]. Supergravity [Des18]. Superheavy [Kra13b, AM12]. surprises [Kar14]. survey [Spi12a]. Swieca [Sch10]. symmetry [Mou13]. synchrotron [Pla12]. systems [GLTZ10, GS12].

table [Kra13b]. Tales [PR18]. technics [DPP^{+}12]. techniques [CM13]. technology [AV24, TWZ19]. telecommunication [MAE^{+}19]. temporal [RD12]. their [AM23, Gal13, Per16, Sch10]. theorem [Bad11, Blu14, BdV22, vN10]. Theoretical [Iof13, MB17a, DB14, dDD^{+}19]. theories [AV24, RV24]. Theory [BGLR17, Ein31, AV24, Bis15, DEN24, Dit15a, Dit16, DM11, DJ13, Eck10, Fic13, FR24, GG19, Haa10a, Ina15, Joh20, Kra11, LJ15, Mas14, OMNM14, RV24, SS17, Sau19, Sch10, Sch11, Som14b, VR17, Wei21, Wu21, Zeh11]. thermodynamic [DD13]. Thermodynamics [MW12, Bor13, LJ15]. thermonuclear [Esc18]. Things [Com29]. thinking [Mou13]. time [GLTZ10, Lev16]. timeline [MB17b]. times [BC18]. today [RV24]. tokamak [Kik18]. tokamaks [PAP21]. told [DFP18]. Tomonaga [Dit16]. torsion [Sch19, Sch20]. Torus [Reb18]. Touschek [BP11]. transfer [AV24, Mas14]. transformations [PP19, PP16]. translation [OM14, OM22, Shi24]. translational [Sch19, Sch20]. transport [LJ15]. traveling [MAE^{+}19]. traveling-wave [MAE^{+}19]. treatise [FK19]. troublesome [Eck10]. tube [MAE^{+}19]. tumultuous [AV24]. turbulence [Eck10, HM18].

Ueber [Shi24]. ultra [KW12]. underlying [LL14]. unified [Sau19]. unique [AV24, Bil13, RMD23]. uniqueness [MPP^{+}12]. universal [Gor24b, Ish17, PB17, Wu21]. universality [Lau24]. universe [Nus14, OM14, OONM17, OM22]. universities [LR20]. University [Ble12, DPP^{+}12, Dor13, GS12, MPP^{+}12, NHK^{+}22, Shl12, Sin12]. upper [Kra13b]. uranium [HS16]. using [FP23, Per16, TWZ19].

vacuum [ABR23]. Varenna [Mar19]. variability [RD12]. Velocity [Bet32, Gen20]. versus [DJ13]. vertical [Sch14a]. Very [LW12]. Very-high [LW12]. view [LPA18]. violation [Cro12]. Virgo [BCC^{+}24]. visibility [BCC^{+}24]. Vlasov [GK23]. vs [Alm20, Bad11, Kre15].

Waiting [Tan16]. walk [Mar19, Nar18]. War [LP24]. Was [vS21]. Wave [LPA18, La 22, Kre11, Kre15, MAE^{+}19, MM24, Ros17, Wu21]. waves [Lor19, Piz16]. Weber [Tri17]. weeks [Des18]. Weinberg [DLL16]. Weiss [RB15]. Weisskopf [Dit15b]. Wheeler [Ric18]. Who [Smi13]. Wick [GS24]. will [Kra11]. Wired [Tri17]. work [FGV17, Shi24]. working [Sau19]. world [AM12, LP24]. wrongs [RD12].

X [Pel12]. X-ray [Pel12].

year [GK23]. Years [Pau13, PAP21, BT21, Haa10b, Iof13, Kar22, Kon20, Mar18, OONM18, Pla12, Spi12a, Wei15]. yesterday [RV24]. Youth [CB15].

Zeitschrift [PMG23, FBB14]. Zocher [SCS22]. Zum [Ein31]. zwischen [PMG23].

References

Agil:2023:SLV

[ABR23] Jonathan Agil, Rémy Battesti, and Carlo Rizzo. On the speed of light in a vacuum in the presence of a magnetic field. European Physical Journal H, 48(1):??, December 2023. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-023-00050-7.

Almeida:2023:AGH

[AJ23] Carla R. Almeida and Maxime J. Jacquet. Analogue gravity and the Hawking effect: historical perspective and literature review. European Physical Journal H, 48(1):??, December 2023. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-023-00063-2.

Ali:2011:JQH

[AK11] A. Ali and G. Kramer. JETS and QCD: a historical review of the discovery of the quark and gluon jets and its impact on QCD. European Physical Journal H, 36(2):245–326, September 2011. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2011-10047-1.

Almeida:2020:SEV

[Alm20] Carla Rodrigues Almeida. Stellar equilibrium vs. gravitational collapse. European Physical Journal H, 45(1):25–48, July 2020. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2019-100045-x;   http://link.springer.com/content/
pdf/10.1140/epjh/e2019-100045-x.pdf.

Armbruster:2012:EPO

[AM12] P. Armbruster and Gottfried Münzenberg. An experimental paradigm opening the world of superheavy elements. European Physical Journal H, 37(2):237–309, July 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2012-20046-7.

Alpher:2023:GGR

[AM23] Victor S. Alpher and Simon A. Mitton. George Gamow and Ralph Alpher: a review of their cosmological collaboration as mentor and protégé 1942–1955. European Physical Journal H, 48(1):??, December 2023. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL https://link.springer.com/article/
10.1140/epjh/s13129-023-00057-0.

Ascuitto:2024:ACC

[AV24] Robert J. Ascuitto and Jan S. Vaagen. The 50th anniversary of the coupled channels Born approximation (CCBA) and the coupled reaction channels (CRC) theories of nucleon transfer reactions (a unique interplay between theory, experiment and computer technology, conducted during the most tumultuous period in modern American society). European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL https://link.springer.com/article/
10.1140/epjh/s13129-023-00060-5.

Badino:2011:MSV

[Bad11] M. Badino. Mechanistic slumber vs. statistical insomnia: the early history of Boltzmann’s H-theorem (1868–1877). European Physical Journal H, 36(3):353–378, November 2011. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2011-10048-5.

Baily:2013:EAM

[Bai13] C. Baily. Early atomic models — from mechanical to quantum (1904–1913). European Physical Journal H, 38(1):1–38, January 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2012-30009-7.

Battimelli:2018:BAP

[BC18] Giovanni Battimelli and Giovanni Ciccotti. Berni Alder and the pioneering times of molecular simulation. European Physical Journal H, 43(3):303–335, August 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.

com/article/10.1140/epjh/e2018-90027-5.

Barneo:2024:APL

[BCC^{+}24] Pablo Barneo, Giuseppe Cabras, Pierre-François Cohadon, Livia Conti, Davide Guerra, Edoardo Milotti, Jerome Novak, Agata Trovato, and Andrea Virtuoso. Addressing the problem of the LIGO–Virgo–KAGRA visibility in the scientific literature. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL

https://link.springer.com/article/10.1140/epjh/s13129-
023-00066-z.

Belyaev:2021:HL

[BCT21] I. Belyaev, G. Carboni, and F. Teubert. The history of LHCb. European Physical Journal H, 46(1):??, December 2021. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-021-00002-z.

Blum:2022:GCT

[BdV22] Alexander S. Blum and Andrés Martínez de Velasco. The genesis of the CPT theorem. European Physical Journal H, 47(1):??, December 2022. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL https://link.springer.com/article/
10.1140/epjh/s13129-022-00037-w.

Beiglbock:2010:E

[Bei10] Wolf Beiglböck. Editorial. European Physical Journal H, 35(1):1–2, July 2010. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467

(electronic). URL http://link.springer.com/content/pdf/10.
1140/epjh/e2010-00005-y.pdf.

Bennett:2010:MII

[Ben10] J. Bennett. MacCullagh’s Ireland: the institutional and cultural space for geometry and physics. European Physical Journal H, 35 (2):123–132, November 2010. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2010-00011-4.

Bertlmann:2020:RRQ

[Ber20] Reinhold A. Bertlmann. Real or not real that is the question . . .. European Physical Journal H, 45(2–3):205–236, September 2020. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/e2020-10022-x.

Bethe:1932:BER

[Bet32] Hans A. Bethe. Bremsformel für Elektronen relativistischer Geschwindigkeit. (German) [Scattering of elections of relativistic velocity]. Zeitschrift für Physik, 76(5–6):293–299, ???? 1932. CO-DEN ZEPYAA. ISSN 0044-3328. See annotation [FBB14].

Bethe:2014:BFE

[Bet14] Hans Bethe. Braking formula for electrons of relativistic speed. European Physical Journal H, 39(5):537–542, December 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/

epjh/e2014-50028-5. See annotation [FBB14].

Beiglbock:2014:E

[BG14] Wolf Beiglböck and Francesco Guerra. Editorial. European Physical Journal H, 39(1):1, February 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.

com/content/pdf/10.1140/epjh/e2014-50005-y.pdf.

Blum:2017:EIS

[BGLR17] Alexander Blum, Domenico Giulini, Roberto Lalli, and Jürgen Renn. Editorial introduction to the special issue “The Renaissance of Einstein’s Theory of Gravitation”. European Physical Journal H, 42(2):95–105, June 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

content/pdf/10.1140/epjh/e2017-80023-3.pdf.

Blum:2017:QGM

[BH17] Alexander Blum and Thiago Hartz. The 1957 quantum gravity meeting in Copenhagen: An analysis of Bryce S. DeWitt’s report. European Physical Journal H, 42(2):107–157, June 2017. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic).

URL http://link.springer.com/content/pdf/10.1140/epjh/
e2017-80015-8.pdf.

Bilenky:2013:NHU

[Bil13] S. M. Bilenky. Neutrino. History of a unique particle. European Physical Journal H, 38(3):345–404, April 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2012-20068-9.

Bistafa:2015:EFF

[Bis15] Sylvio R. Bistafa. Euler’s friction of fluids theory and the estimation of fountain jet heights. European Physical Journal H, 40 (3):375–384, September 2015. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2015-60031-2.

Blum:2024:E

[BL24] Alexander S. Blum and Matteo Leone. Editorial. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.

springer.com/article/10.1140/epjh/s13129-024-00081-8.

Bleher:2012:SMS

[Ble12] Pavel Bleher. From the seminar on Mathematical Statistical Physics in Moscow State University, 1962–1994. Hierarchical models and renormalisation group critical phenomena in the Dyson hierarchical model and renormalisation group. European Physical Journal H, 37 (4):605–618, September 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2012-10053-x.

Battimelli:2020:MSI

[BLR20] Giovanni Battimelli, Adele La Rana, and Paolo Rossi. Masters and students in Italian physics between the 19th and 20th centuries: the Felici–Bartoli–Stracciati–Corbino case. European Physical Journal H, 45(2–3):107–121, September 2020. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.

springer.com/article/10.1140/epjh/e2020-10016-y.

Blum:2014:NPG

[Blu14] Alexander Blum. From the necessary to the possible: the genesis of the spin–statistics theorem. European Physical Journal H, 39 (5):543–574, December 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2014-50022-5.

Bonolis:2017:SSC

[Bon17] Luisa Bonolis. Stellar structure and compact objects before 1940: Towards relativistic astrophysics. European Physical Journal H, 42(2):311–393, June 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

content/pdf/10.1140/epjh/e2017-80014-4.pdf.

Bordoni:2013:RTA

[Bor13] Stefano Bordoni. Routes towards an abstract thermodynamics in the late nineteenth century. European Physical Journal H, 38(5): 617–660, December 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2013-40028-7.

Borrelli:2015:SHB

[Bor15] Arianna Borrelli. The story of the Higgs boson: the origin of mass in early particle physics. European Physical Journal H, 40(1):1–52, March 2015. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2014-50026-9.

Boughn:2013:FHM

[Bou13] Stephen Boughn. Fritz Hasenöhrl and E = mc^{2}. European Physical Journal H, 38(2):261–278, March 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2012-30061-5.

Bonolis:2011:BTP

[BP11] Luisa Bonolis and Giulia Pancheri. Bruno Touschek: particle physicist and father of the e^{+} e^{−} collider. European Physical Journal H, 36(1):1–61, July 2011. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2011-10044-1.

Boughn:2013:PAM

[BR13] Stephen Boughn and Marcel Reginatto. A pedestrian approach to the measurement problem in quantum mechanics. European Physical Journal H, 38(4):443–470, September 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2013-40004-9.

Baksalary:2021:MPI

[BT21] Oskar Maria Baksalary and Götz Trenkler. The Moore–Penrose inverse: a hundred years on a frontline of physics research. European Physical Journal H, 46(1):??, December 2021. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-021-00011-y.

Bethke:2022:JEP

[BW22] S. Bethke and A. Wagner. The JADE experiment at the PETRA e^{+}e^{−} collider: history, achievements and revival. European Physical Journal H, 47(1):??, December 2022. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.

springer.com/article/10.1140/epjh/s13129-022-00047-8.

Cenadelli:2015:YPE

[CB15] Davide Cenadelli and Andrea Bernagozzi. Youth plus experience: the discovery of 51 Pegasi b. European Physical Journal H, 40(4– 5):527–552, December 2015. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2015-60041-5.

Carlson:2011:NIS

[CD11] Per Carlson and Alessandro De Angelis. Nationalism and internationalism in science: the case of the discovery of cosmic rays. European Physical Journal H, 35(4):309–329, April 2011. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2011-10033-6.

Casalbuoni:2022:BHF

[CDM22] Roberto Casalbuoni, Daniele Dominici, and Massimo Mazzoni. A brief history of Florentine physics from the 1920s to the end of the 1960s. European Physical Journal H, 47(1):??, December 2022. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-022-00048-7.

Costantini:2024:PJM

[CF24a] Domenico Costantini and Carlo Ferigato. Pascual Jordan: from matrix multiplication to interference law. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.springer.

com/article/10.1140/epjh/s13129-024-00073-8. See publisher

correction [CF24b].

Costantini:2024:PCP

[CF24b] Domenico Costantini and Carlo Ferigato. Publisher correction: Pascual Jordan: from matrix multiplication to interference law. European Physical Journal H, 49(1):??, December 2024.

CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-024-00077-4. See [CF24a].

Caspers:2012:HSB

[CM12] F. Caspers and D. Möhl. History of stochastic beam cooling and its application in many different projects. European Physical Journal H, 36(4):601–632, March 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2012-20037-8.

Cremer:2013:RET

[CM13] Christoph Cremer and Barry R. Masters. Resolution enhancement techniques in microscopy. European Physical Journal H, 38(3):281–344, April 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

content/pdf/10.1140/epjh/e2012-20060-1.pdf.

Compton:1929:WTMa

[Com29] Arthur H. Compton. What things are made of — I. Scientific American, 140(2):110–113, February 1929. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:/

/www.nature.com/scientificamerican/journal/v140/n2/pdf/
scientificamerican0229-110.pdf.

Craik:2012:LKF

[Cra12] Alex D. D. Craik. Lord Kelvin on fluid mechanics. European Physical Journal H, 37(1):75–114, June 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2012-30004-y.

Cronin:2011:CRC

[Cro11] J. W. Cronin. The 1953 Cosmic Ray Conference at Bagnères de Bigorre: the birth of sub atomic physics. European Physical Journal H, 36(2):183–201, September 2011. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.

com/article/10.1140/epjh/e2011-20014-4.

Cronin:2012:DCV

[Cro12] J. W. Cronin. The discovery of CP violation. European Physical Journal H, 36(4):487–508, March 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2011-20057-4.

deArrieta:2022:BIC

[dA22] Iñigo González de Arrieta. Beyond the infrared: a centenary of Heinrich Rubens’s death. European Physical Journal H, 47(1):??, December 2022. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL https://link.springer.com/article/
10.1140/epjh/s13129-022-00044-x.

Dahmen:2015:PAR

[Dah15] Sílvio R. Dahmen. On pendulums and air resistance. European Physical Journal H, 40(3):337–373, September 2015. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2015-50054-8.

Darrigol:2010:JME

[Dar10] Olivier Darrigol. James MacCullagh’s ether: An optical route to Maxwell’s equations? European Physical Journal H, 35(2):133–172, November 2010. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2010-00009-3.

Darrigol:2019:MFQ

[Dar19] Olivier Darrigol. The magic of Feynman’s QED: from field-less electrodynamics to the Feynman diagrams. European Physical Journal H, 44(4–5):349–369, November 2019. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.

springer.com/article/10.1140/epjh/e2019-100025-2.

Darrigol:2023:HRB

[Dar23] Olivier Darrigol. A history of the relation between fluctuation and dissipation. European Physical Journal H, 48(1):??, December 2023. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-023-00058-z.

DiCastro:2014:BTC

[DB14] Carlo Di Castro and Luisa Bonolis. The beginnings of theoretical condensed matter physics in Rome: a personal remembrance. European Physical Journal H, 39(1):3–36, February 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2013-40043-5.

Deville:2013:CLB

[DD13] Alain Deville and Yannick Deville. Clarifying the link between von Neumann and thermodynamic entropies. European Physical Journal H, 38(1):57–81, January 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.

com/article/10.1140/epjh/e2012-30032-0.

dAlessandro:2019:SRI

[dDD^{+}19] Vincenzo d’Alessandro, Santolo Daliento, Marco Di Mauro, Salvatore Esposito, and Adele Naddeo. Searching for a response: the intriguing mystery of Feynman’s theoretical reference amplifier. European Physical Journal H, 44(4–5):331–347, November 2019. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2019-90071-6.

DiMauro:2024:DCE

[DEN24] Marco Di Mauro, Salvatore Esposito, and Adele Naddeo. The development of the concept of exchange forces in the 1930s: close encounters between Europe and Japan and the birth of nuclear theory. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-024-00078-3.

Deser:2018:BHG

[Des18] S. Deser. A brief history (and geography) of supergravity: the first 3 weeks … and after. European Physical Journal H, 43(3):281–291, August 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2018-90005-3.

DeWitt:2017:ERS

[DeW17] Bryce S. DeWitt. Exploratory research session on the quantization of the gravitational field. European Physical Journal H, 42(2):159–176, June 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

content/pdf/10.1140/epjh/e2017-80016-0.pdf.

Diamond:2018:EIS

[DFP18] Patrick H. Diamond, Uriel Frisch, and Yves Pomeau. Editorial introduction to the special issue “Plasma physics in the 20th century as told by players”. European Physical Journal H, 43(4–5):337–353, December 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/
10.1140/epjh/e2018-90061-5;   http://link.springer.com/
content/pdf/10.1140/epjh%2Fe2018-90061-5.pdf.

Dittrich:2015:CQF

[Dit15a] Walter Dittrich. The cofounder of quantum field theory: Pascual Jordan. European Physical Journal H, 40(2):241–260, March 2015. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2015-60005-9.

Dittrich:2015:PWA

[Dit15b] Walter Dittrich. On the Pauli–Weisskopf anti-Dirac paper. European Physical Journal H, 40(2):261–278, March 2015. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2015-60006-1.

Dittrich:2016:TTS

[Dit16] Walter Dittrich. On Tomonaga’s theory of split-anode magnetrons. European Physical Journal H, 41(2):165–180, June 2016. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2016-70005-7.

Duncan:2013:NMY

[DJ13] A. Duncan and M. Janssen. (never) mind your p’s and q’s: von Neumann versus Jordan on the foundations of quantum theory. European Physical Journal H, 38(2):175–259, March 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2012-30024-5.

Dore:2016:MWA

[DLL16] Ubaldo Dore, Pier Ferruccio Loverre, and Lucio Ludovici. Measurement of the Weinberg angle in neutrino interactions. European Physical Journal H, 41(2):137–155, June 2016. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2016-70006-y.

Dore:2019:HAN

[DLL19] Ubaldo Dore, Pier Loverre, and Lucio Ludovici. History of accelerator neutrino beams. European Physical Journal H, 44(4–5):271–305, November 2019. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2019-90032-x.

Dosch:2011:FRQ

[DM11] H. G. Dosch and V. F. Müller. The facets of relativistic quantum field theory. European Physical Journal H, 35(4):331–375, April 2011. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2011-10030-6.

Dore:2012:PCP

[Dor12] U. Dore. Physics with charm particles produced in neutrino interactions. A historical recollection. European Physical Journal H, 37(1):115–137, June 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2012-20019-x.

Dore:2013:NEP

[Dor13] Ubaldo Dore. Neutrino experiments in the Physics Department of Rome ‘Sapienza’ University. European Physical Journal H, 38 (5):703–712, December 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2013-40010-8.

Dinaburg:2012:SMS

[DPP^{+}12] E. Dinaburg, E. A. Pechersky, S. A. Pirogov, S. Shlosman, and Yu. M. Suhov. From the seminar on Mathematical Statistical Physics in Moscow State University, 1962–1994. Contour technics. European Physical Journal H, 37(4):619–637, September 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2012-10052-6.

Darrigol:2010:HJM

[DS10] Olivier Darrigol and Samson Shatashvili. In honor of James Mac-Cullagh (1809–1847). European Physical Journal H, 35(2):111, November 2010. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/content/
pdf/10.1140/epjh/e2010-00010-6.pdf.

Duarte:2012:OQE

[Dua12] F. J. Duarte. The origin of quantum entanglement experiments based on polarization measurements. European Physical Journal H, 37(2):311–318, July 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2012-20047-y.

Esposito:2024:PLS

[EA24] Salvatore Esposito and Alessandro Amabile. Percolating lives: statistical mechanics in Naples. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL https://link.springer.com/article/
10.1140/epjh/s13129-024-00076-5.

Eckert:2010:TBH

[Eck10] M. Eckert. The troublesome birth of hydrodynamic stability theory: Sommerfeld and the turbulence problem. European Physical Journal H, 35(1):29–51, July 2010. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2010-00003-3.

Eckhardt:2012:HDS

[Eck12] Gisela Eckhardt. History of the discovery of the stimulated Raman effect at the Hughes Research Laboratories. European Physical Journal H, 37(5):793–796, October 2012. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic).

URL http://link.springer.com/content/pdf/10.1140/epjh/
e2012-30025-9.pdf.

Eckert:2014:HSE

[Eck14] Michael Eckert. How Sommerfeld extended Bohr’s model of the atom (1913–1916). European Physical Journal H, 39(2):141–156, April 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2013-40052-4.

Einstein:1931:KPA

[Ein31] Albert Einstein. Zum Kosmologischen Problem der allgemeinen Relativitätstheorie. (German) [On the cosmological problem of the General Theory of Relativity]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):235–237, 1931.

Ellis:2014:SHA

[Ell14] George F. R. Ellis. Stephen Hawking’s 1966 Adams Prize Essay. European Physical Journal H, 39(4):403–411, November 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2014-50014-x.

Escande:2018:TFH

[Esc18] D. F. Escande. From thermonuclear fusion to Hamiltonian chaos. European Physical Journal H, 43(4–5):397–420, December 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2016-70063-5.

Fontes:2014:AHB

[FBB14] Christopher J. Fontes, Christopher J. Bostock, and Klaus Bartschat. Annotation of Hans Bethe’s paper, Zeitschrift für Physik 76, 293 (1932), “Braking Formula for Electrons of Relativistic Speed”. European Physical Journal H, 39(5):517–536, December 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2014-50027-1. See [Bet32, Bet14].

Frisch:2017:CLH

[FGV17] Uriel Frisch, Gérard Grimberg, and Barbara Villone. A contemporary look at Hermann Hankel’s 1861 pioneering work on Lagrangian fluid dynamics. European Physical Journal H, 42(4–5):537–556, December 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic).

Folk:2022:SFS

[FH22] Reinhard Folk and Yurij Holovatch. Schottky’s forgotten step to the Ising model. European Physical Journal H, 47(1):??, December 2022. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-022-00041-0.

Fick:2013:BHA

[Fic13] D. Fick. Bothe’s 1925 heuristic assumption in the dawn of quantum field theory. European Physical Journal H, 38(1):39–55, January 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2012-20023-1.

Fischbach:2015:FFP

[Fis15] Ephraim Fischbach. The fifth force: A personal history. European Physical Journal H, 40(4–5):385–467, December 2015. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2015-60044-5.

Frisvad:2019:LLH

[FK19] Jeppe Revall Frisvad and Helge Kragh. On Ludvig Lorenz and his 1890 treatise on light scattering by spheres. European Physical Journal H, 44(2):137–160, August 2019. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2019-100022-y.

Fraser:2023:NEH

[FP23] Doreen Fraser and Maria Papageorgiou. Note on episodes in the history of modeling measurements in local spacetime regions using QFT. European Physical Journal H, 48(1):??, December 2023. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-023-00064-1.

Fraser:2024:PEF

[FR24] James D. Fraser and Kasia Rejzner. Perturbative expansions and the foundations of quantum field theory. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.springer.

com/article/10.1140/epjh/s13129-024-00075-6.

Franklin:2013:MMP

[Fra13] Allan Franklin. Millikan’s measurement of Planck’s constant. European Physical Journal H, 38(5):573–594, December 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2013-40021-3.

Fredenhagen:2010:LBC

[Fre10] Klaus Fredenhagen. Lille 1957: The birth of the concept of local algebras of observables. European Physical Journal H, 35(3):239–241, November 2010. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2010-10040-y.

Friedman:2012:PIP

[Fri12] J. I. Friedman. Peering inside the proton. European Physical Journal H, 36(4):469–485, March 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.

com/article/10.1140/epjh/e2012-30002-3.

Frisch:2014:CAF

[FV14] Uriel Frisch and Barbara Villone. Cauchy’s almost forgotten Lagrangian formulation of the Euler equation for 3D incompressible flow. European Physical Journal H, 39(3):325–351, September 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2014-50016-6.

Gallavotti:2013:ALM

[Gal13] Giovanni Gallavotti. Aspects of Lagrange’s Mechanics and their legacy. European Physical Journal H, 38(5):595–615, December 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2013-40029-y.

Gallavotti:2016:EHP

[Gal16] Giovanni Gallavotti. Ergodicity: a historical perspective. equilibrium and nonequilibrium. European Physical Journal H, 41(3): 181–259, September 2016. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

accesspage/article/10.1140/epjh/e2016-70030-8.

Genthon:2020:CVH

[Gen20] Arthur Genthon. The concept of velocity in the history of Brownian motion. European Physical Journal H, 45(1):49–105, July 2020. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2020-10009-8.

Galvagno:2019:LSC

[GG19] Mariano Galvagno and Gaston Giribet. Luis Santaló and classical field theory. European Physical Journal H, 44(4–5):381–389, November 2019. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2019-100038-9.

Gabovich:2023:AVH

[GK23] Alexander M. Gabovich and Vladimir I. Kuznetsov. Anatoly Vlasov heritage: 60-year-old controversy. European Physical Journal H, 48(1):??, December 2023. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.springer.

com/article/10.1140/epjh/s13129-023-00051-6.

Goldstein:2010:LTB

[GLTZ10] S. Goldstein, J. L. Lebowitz, R. Tumulka, and N. Zanghì. Longtime behavior of macroscopic quantum systems. European Physical Journal H, 35(2):173–200, November 2010. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2010-00007-7.

Goenner:2017:GRG

[Goe17] Hubert Goenner. General relativity and the growth of a subdiscipline “gravitation” in Germany. European Physical Journal H, 42(3):395–430, August 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic).

Gorelik:2024:DIH

[Gor24a] Gennady Gorelik. The drama of ideas in the history of quantum gravity: Niels Bohr, Lev Landau, and Matvei Bronstein. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-024-00080-9.

Gorelik:2024:HFA

[Gor24b] Gennady Gorelik. How a falling apple could have helped Newton discover universal gravity. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL https://link.springer.com/article/
10.1140/epjh/s13129-023-00065-0.

Gurevich:2012:SMS

[GS12] B. M. Gurevich and Y. M. Suhov. From the seminar on Mathematical Statistical Physics in Moscow State University, 1962–1994. Dynamical systems of infinitely many particles. European Physical Journal H, 37(4):639–658, September 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2012-10054-2.

Gould:2022:FFF

[GS22] Christopher R. Gould and Eduard I. Sharapov. Fermi’s favorite figure: the history of the pseudopotential concept in atomic physics and neutron physics. European Physical Journal H, 47(1):??, December 2022. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL https://link.springer.com/article/
10.1140/epjh/s13129-022-00042-z.

Gould:2024:GCW

[GS24] Christopher R. Gould and Eduard I. Sharapov. Gian-Carlo Wick and neutron physics in the 1930s. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.springer.

com/article/10.1140/epjh/s13129-024-00072-9.

Haag:2010:DAA

[Haa10a] Rudolf Haag. Discussion of the ‘axioms’ and the asymptotic properties of a local field theory with composite particles. European Physical Journal H, 35(3):243–253, November 2010. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2010-10041-3.

Haag:2010:LAL

[Haa10b] Rudolf Haag. Local algebras. A look back at the early years and at some achievements and missed opportunities. European Physical Journal H, 35(3):255–261, November 2010. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2010-10042-7.

Haag:2010:SPS

[Haa10c] Rudolf Haag. Some people and some problems met in half a century of commitment to mathematical physics. European Physical Journal H, 35(3):263–307, November 2010. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2010-10032-4.

Hales:2024:RMJ

[Hal24] Thomas Hales. Robert Millikan, Japanese internment, and eugenics. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-024-00068-5.

Hawking:2014:SGS

[Haw14] Stephen Hawking. Singularities and the geometry of spacetime. European Physical Journal H, 39(4):413–503, November 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2014-50013-6.

Helbig:2021:AAF

[Hel21] Phillip Helbig. Arguments against the flatness problem in classical cosmology: a review. European Physical Journal H, 46(1):??, December 2021. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL https://link.springer.com/article/
10.1140/epjh/s13129-021-00006-9.

Hicks:2012:CP

[Hic12] Kenneth H. Hicks. On the conundrum of the pentaquark. European Physical Journal H, 37(1):1–31, June 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2012-20032-0.

Hasegawa:2018:STS

[HM18] Akira Hasegawa and Kunioki Mima. Strong turbulence, selforganization and plasma confinement. European Physical Journal H, 43(4–5):499–521, December 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.

springer.com/article/10.1140/epjh/e2018-90033-4.

Harlander:2024:DCM

[HM24] Robert V. Harlander and Jean-Philippe Martinez. The development of computational methods for Feynman diagrams. European Physical Journal H, 49(1):4:1–4:45, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-024-00067-6.

Harlander:2023:CEP

[HMS23a] Robert Harlander, Jean-Philippe Martinez, and Gregor Schiemann. Correction: The end of the particle era? European Physical Journal H, 48(1):??, December 2023. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL

https://link.springer.com/article/10.1140/epjh/s13129-

023-00059-y. See [HMS23b].

Harlander:2023:EPE

[HMS23b] Robert Harlander, Jean-Philippe Martinez, and Gregor Schiemann. The end of the particle era? European Physical Journal H, 48(1):??, December 2023. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL https://link.springer.com/article/
10.1140/epjh/s13129-023-00053-4. See correction [HMS23a].

Holian:2021:EBB

[Hol21] Brad Lee Holian. Exploring the boundary between atoms and the continuum by computers: a personal history. European Physical Journal H, 46(1):??, December 2021. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.

springer.com/article/10.1140/epjh/s13129-021-00010-z.

Heilbron:2023:MMM

[HR23] John L. Heilbron and Carlo Rovelli. Matrix mechanics mis-prized: Max Born’s belated nobelization. European Physical Journal H, 48(1):??, December 2023. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://arxiv.org/

abs/2306.00842;   https://link.springer.com/article/10.
1140/epjh/s13129-023-00056-1.

Hahn:2016:PRI

[HS16] O. Hahn and F. Strassmann. On the production of radium isotopes from uranium caused by irradiation with fast and decelerated neutrons. European Physical Journal H, 41(3):261–264, September 2016. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/accesspage/article/
10.1140/epjh/e2016-70025-1.

Hubner:2012:CIS

[Hüb12] Kurt Hübner. The CERN Intersecting Storage Rings (ISR). European Physical Journal H, 36(4):509–522, March 2012. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2011-20058-8.

Imre:2016:FKL

[IN16] Attila R. Imre and Josef Novotný. Fractals and the Korcak-law: a history and a correction. European Physical Journal H, 41(1):69– 91, April 2016. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2016-60039-8.

Inaba:2015:DET

[Ina15] Hajime Inaba. The development of ensemble theory. European Physical Journal H, 40(4–5):489–526, December 2015. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2015-60034-2.

Ioffe:2013:FDY

[Iof13] B. L. Ioffe. The first dozen years of the history of ITEP Theoretical Physics Laboratory. European Physical Journal H, 38(1):83–135, January 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2012-30008-3.

Ishiwara:2017:UMQ

[Ish17] Jun Ishiwara. The universal meaning of the quantum of action. European Physical Journal H, 42(4–5):523–536, December 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/

epjh/e2017-80041-1. See analysis [PB17].

Jentschura:2014:ADF

[JN14] U. D. Jentschura and I. Nándori. Attempts at a determination of the fine-structure constant from first principles: a brief historical overview. European Physical Journal H, 39(5):591–613, December 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2014-50044-7.

Johnson:2020:CGD

[Joh20] Benjamin Johnson. Charles Galton Darwin’s 1922 quantum theory of optical dispersion. European Physical Journal H, 45(1):1–23, July 2020. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/
10.1140/epjh/e2020-80058-7;   http://link.springer.com/
content/pdf/10.1140/epjh/e2020-80058-7.pdf.

Karlsson:2014:PMI

[Kar14] Erik B. Karlsson. The positive muon implanted in metals — a story full of surprises. European Physical Journal H, 39(3):303–323, September 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2014-50018-2.

Karlsson:2022:IDC

[Kar22] Erik B. Karlsson. Internal dynamics in condensed matter, as studied by spin relaxation: some examples from 75 years. European Physical Journal H, 47(1):??, December 2022. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL

https://link.springer.com/article/10.1140/epjh/s13129-
021-00030-9.

Kennefick:2017:BPQ

[Ken17] Daniel Kennefick. The binary pulsar and the quadrupole formula controversy. European Physical Journal H, 42(2):293–310, June 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467

(electronic). URL http://link.springer.com/content/pdf/10.
1140/epjh/e2016-70059-2.pdf.

Kikuchi:2018:LTJ

[Kik18] Mitsuru Kikuchi. The large tokamak JT-60: a history of the fight to achieve the Japanese fusion research mission. European Physical Journal H, 43(4–5):551–577, December 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2018-90054-2.

Kajita:2012:OKE

[KKS12] T. Kajita, M. Koshiba, and A. Suzuki. On the origin of the Kamiokande experiment and neutrino astrophysics. European Physical Journal H, 37(1):33–73, June 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.

springer.com/article/10.1140/epjh/e2012-30007-y.

Kono:2020:RKH

[Kon20] Hiroto Kono. Ryogo Kubo in his formative years as a physicist. European Physical Journal H, 45(2–3):175–204, September 2020. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/e2020-10003-8.

Kragh:2011:COB

[Kra11] Helge Kragh. Conceptual objections to the Bohr atomic theory — do electrons have a “free will”? European Physical Journal H, 36 (3):327–352, November 2011. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2011-20031-x.

Kragh:2013:NCB

[Kra13a] Helge Kragh. Nordic cosmogonies: Birkeland, Arrhenius and finde-siècle cosmical physics. European Physical Journal H, 38(4): 549–572, September 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2013-40014-0.

Kragh:2013:SEU

[Kra13b] Helge Kragh. Superheavy elements and the upper limit of the periodic table: early speculations. European Physical Journal H, 38(3):411–431, April 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2012-30043-7.

Kragh:2014:NPP

[Kra14] Helge Kragh. The names of physics: plasma, fission, photon. European Physical Journal H, 39(3):263–281, September 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2014-50007-7.

Kragh:2016:SSE

[Kra16] Helge Kragh. The source of solar energy, ca. 1840–1910: From meteoric hypothesis to radioactive speculations. European Physical Journal H, 41(4–5):365–394, November 2016. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/accesspage/article/
10.1140/epjh/e2016-70045-7.

Krehl:2011:SWP

[Kre11] Peter O. K. Krehl. Shock wave physics and detonation physics — a stimulus for the emergence of numerous new branches in science and engineering. European Physical Journal H, 36(1):85–152, July 2011. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2011-10037-x.

Krehl:2015:CRH

[Kre15] Peter O. K. Krehl. The classical Rankine–Hugoniot jump conditions, an important cornerstone of modern shock wave physics: ideal assumptions vs. reality. European Physical Journal H, 40 (2):159–204, March 2015. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2015-50010-4.

Kampert:2012:EAS

[KW12] Karl-Heinz Kampert and Alan A. Watson. Extensive air showers and ultra high-energy cosmic rays: a historical review. European Physical Journal H, 37(3):359–412, August 2012. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2012-30013-x.

Rana:2022:EFA

[La 22] Adele La Rana. EUROGRAV 1986–1989: the first attempts for a European Interferometric Gravitational Wave Observatory. European Physical Journal H, 47(1):??, December 2022. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-022-00036-x.

Lange:2014:LI

[Lan14] Ludwig Lange. On the Law of Inertia. European Physical Journal H, 39(2):251–262, April 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2013-40040-5.

Laudisa:2024:BNU

[Lau24] Federico Laudisa. Bohr and von Neumann on the universality of quantum mechanics: materials for the history of the quantum measurement process. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL https://link.springer.com/article/
10.1140/epjh/s13129-024-00082-7.

Lebowitz:2017:LSM

[LB17] Joel L. Lebowitz and Luisa Bonolis. A life in statistical mechanics. European Physical Journal H, 42(1):1–21, April 2017. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic).

URL http://link.springer.com/content/pdf/10.1140/epjh/
e2017-80006-9.pdf.

Lessel:2024:HPH

[Les24] Bernadette Lessel. From history of physics to “history for physics”. European Physical Journal H, 49(1):??, December 2024.

CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-024-00084-5.

Levine:2016:HTF

[Lev16] Judah Levine. The history of time and frequency from antiquity to the present day. European Physical Journal H, 41(1):1– 67, April 2016. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2016-70004-3.

Levesque:2019:OCS

[LH19] D. Levesque and J. P. Hansen. The origin of computational statistical mechanics in France. European Physical Journal H, 44(1):37–46, February 2019. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2018-90041-y.

Lebon:2015:EHE

[LJ15] G. Lebon and D. Jou. Early history of extended irreversible thermodynamics (1953–1983): An exploration beyond local equilibrium and classical transport theory. European Physical Journal H, 40 (2):205–240, March 2015. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2014-50033-0.

Labuda:2014:MUD

[LL14] Cecille Labuda and Iwo Labuda. On the mathematics underlying dispersion relations. European Physical Journal H, 39(5):575–589, December 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2014-50021-1.

Lorenz:2019:LPO

[Lor19] Ludvig Lorenz. Light propagation in and outside a sphere illuminated by plane waves of light. European Physical Journal H, 44(2):77–135, August 2019. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2019-100021-6. English translation by

Jeppe Revall Frisvad of original Danish article Lysbevægelser i og uden for en af plane Lysbølger belyst Kugle in Det kongelige danske Videnskabernes Selskabs Skrifter 6(6): 1–62, 1890.

Lalli:2024:BSF

[LP24] Roberto Lalli and Paolo Politi. The birth of StatPhys: the 1949 Florence conference at the juncture of national and international physics reconstruction after World War II. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.

springer.com/article/10.1140/epjh/s13129-024-00079-2.

Laval:2018:WPW

[LPA18] Guy Laval, Denis Pesme, and Jean-Claude Adam. Wave–particle and wave–wave interactions in hot plasmas: a French historical point of view. European Physical Journal H, 43(4–5):421–458, December 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2016-70050-2.

LaRana:2020:BQM

[LR20] Adele La Rana and Paolo Rossi. The blossoming of quantum mechanics in Italy: the roots, the context and the first spreading in Italian universities (1900–1947). European Physical Journal H, 45 (4–5):237–257, December 2020. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.springer.

com/article/10.1140/epjh/e2020-10044-0.

Lorenz:2012:VHE

[LW12] E. Lorenz and R. Wagner. Very-high energy gamma-ray astronomy. European Physical Journal H, 37(3):459–513, August 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2012-30016-x.

Minenna:2019:TWT

[MAE^{+}19] Damien F. G. Minenna, Frédéric André, Yves Elskens, Jean-François Auboin, Fabrice Doveil, Jérôme Puech, and Élise Duverdier. The traveling-wave tube in the history of telecommunication. European Physical Journal H, 44(1):1–36, February 2019. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2018-90023-1.

Mareschal:2018:EYC

[Mar18] Michel Mareschal. Early years of computational statistical mechanics. European Physical Journal H, 43(3):293–302, August 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2018-90006-7.

Mareschal:2019:VCK

[Mar19] Michel Mareschal. From Varenna (1970) to Como (1995): Kurt Binder’s long walk in the land of criticality. European Physical Journal H, 44(2):161–179, August 2019. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.

com/article/10.1140/epjh/e2019-100016-3.

Masters:2014:PFR

[Mas14] B. R. Masters. Paths to Förster’s resonance energy transfer (FRET) theory. European Physical Journal H, 39(1):87–139, February 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2013-40007-9.

Maiani:2017:CTP

[MB17a] Luciano Maiani and Luisa Bonolis. The charm of theoretical physics (1958–1993). European Physical Journal H, 42(4–5):611–661, December 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/content/
pdf/10.1140/epjh/e2017-80040-9.pdf.

Maiani:2017:LTP

[MB17b] Luciano Maiani and Luisa Bonolis. The LHC timeline: a personal recollection (1980–2012). European Physical Journal H, 42(4–5): 475–505, December 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

content/pdf/10.1140/epjh/e2017-80052-8.pdf.

Mayor:2018:EBN

[MC18] Michel Mayor and Davide Cenadelli. Exoplanets — the beginning of a new era in astrophysics. European Physical Journal H, 43(1): 1–41, April 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2018-80063-1.

Minin:2024:BGB

[MM24] Igor V. Minin and Oleg V. Minin. The breakup of gas bubbles by a shock wave: brief historical background. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.

springer.com/article/10.1140/epjh/s13129-024-00071-w.

Mouchet:2013:RFF

[Mou13] Amaury Mouchet. Reflections on the four facets of symmetry: how physics exemplifies rational thinking. European Physical Journal H, 38(5):661–702, December 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2013-40018-4.

Montgomery:2013:LBP

[MP13] H. E. Montgomery, Jr. and V. I. Pupyshev. On lower bounds for polarisability. European Physical Journal H, 38(4):519–534, September 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2013-30056-8.

Minlos:2012:SMS

[MPP^{+}12] R. A. Minlos, E. A. Pechersky, S. A. Pirogov, S. Shlosman, and Yu. M. Suhov. From the seminar on Mathematical Statistical Physics in Moscow State University, 1962–1994. Gibbs random fields on the lattice. Definitions, existence, uniqueness. European Physical Journal H, 37(4):571–594, September 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2012-10049-7.

Macuglia:2022:EPD

[MRC22] Daniele Macuglia, Benoît Roux, and Giovanni Ciccotti. The emergence of protein dynamics simulations: how computational statistical mechanics met biochemistry. European Physical Journal H, 47(1):??, December 2022. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.springer.

com/article/10.1140/epjh/s13129-022-00043-y.

Muller:2012:DOG

[Mül12] Dietrich Müller. Direct observations of galactic cosmic rays. European Physical Journal H, 37(3):413–458, August 2012. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2012-30017-2.

Monfort-Urkizu:2023:WN

[MUN23] Beñat Monfort-Urkizu and Jaume Navarro. What’s in a name? European Physical Journal H, 48(1):??, December 2023. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-023-00054-3.

Muller:2012:TIP

[MW12] Ingo Müller and Wolf Weiss. Thermodynamics of irreversible processes — past and present. European Physical Journal H, 37(2):139– 236, July 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2012-20029-1.

Myers:2012:LO

[Mye12] Steve Myers. LEP operation. European Physical Journal H, 36 (4):563–577, March 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2011-20062-y.

Narlikar:2018:EMC

[Nar18] Jayant V. Narlikar. The evolution of modern cosmology as seen through a personal walk across six decades. European Physical Journal H, 43(1):43–72, April 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.

com/article/10.1140/epjh/e2017-80048-5.

Novosyadlyj:2022:SKD

[NHK^{+}22] B. Novosyadlyj, B. Hnatyk, Yu. Kulinich, B. Melekh, O. Petruk, R. Plyatsko, M. Tsizh, M. Vavrukh, and N. Virun. Samuil Kaplan and the development of astrophysical research at the Lviv University. European Physical Journal H, 47(1):??, December 2022. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-022-00045-w.

Nielsen:2022:OGP

[Nie22] N. K. Nielsen. On the origin of the Gullstrand–Painlevé coordinates. European Physical Journal H, 47(1):??, December 2022. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-022-00038-9.

Novosyadlyj:2018:C

[Nov18] Bohdan Novosyadlyj. Century of Λ. European Physical Journal H, 43(3):267–280, August 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2018-90007-y.

Nussbaumer:2014:ECH

[Nus14] Harry Nussbaumer. Einstein’s conversion from his static to an expanding universe. European Physical Journal H, 39(1):37–62, February 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2013-40037-6.

Oldofredi:2023:ODE

[Old23] Andrea Oldofredi. Orthodox or dissident? The evolution of Bohm’s ontological reflections in the 1950s. European Physical Journal H, 48(1):??, December 2023. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.springer.

com/article/10.1140/epjh/s13129-023-00062-3.

ORaifeartaigh:2014:ECM

[OM14] C. O’Raifeartaigh and B. McCann. Einstein’s cosmic model of 1931 revisited: an analysis and translation of a forgotten model of the universe. European Physical Journal H, 39(1):63–85, February 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.

1140/epjh/e2013-40038-x. Contains English translation on pages 82–85 of [Ein31]. See correction [OM22].

ORaifeartaigh:2022:CEC

[OM22] C. O’Raifeartaigh and B. McCann. Correction to: Einstein’s cosmic model of 1931 revisited: an analysis and translation of a forgotten model of the universe. European Physical Journal H, 47(1):??, December 2022. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL https://link.springer.com/article/
10.1140/epjh/s13129-022-00040-1. See [OM14].

ORaifeartaigh:2014:ESS

[OMNM14] Cormac O’Raifeartaigh, Brendan McCann, Werner Nahm, and Simon Mitton. Einstein’s steady-state theory: an abandoned model of the cosmos. European Physical Journal H, 39(3):353–367, September 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2014-50011-x.

ORaifeartaigh:2021:HPR

[OOM21] Cormac O’Raifeartaigh, Michael O’Keeffe, and Simon Mitton. Historical and philosophical reflections on the Einstein–de Sitter model. European Physical Journal H, 46(1):??, December 2021. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-021-00007-8.

ORaifeartaigh:2015:ECR

[OONM15] Cormac O’Raifeartaigh, Michael O’Keeffe, Werner Nahm, and Simon Mitton. Einstein’s cosmology review of 1933: a new perspective on the Einstein–de Sitter model of the cosmos. European Physical Journal H, 40(3):301–335, September 2015. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2015-50061-y.

ORaifeartaigh:2017:ESM

[OONM17] Cormac O’Raifeartaigh, Michael O’Keeffe, Werner Nahm, and Simon Mitton. Einstein’s 1917 static model of the universe: a centennial review. European Physical Journal H, 42(3):431–474, August 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic).

ORaifeartaigh:2018:OHY

[OONM18] Cormac O’Raifeartaigh, Michael O’Keeffe, Werner Nahm, and Simon Mitton. One hundred years of the cosmological constant: from ‘superfluous stunt” to dark energy. European Physical Journal H, 43(1):73–117, April 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2017-80061-7.

Petrov:2021:YNP

[PAP21] M. P. Petrov, V. I. Afanasyev, and S. Ya. Petrov. 60 years of neutral particle analysis: from early tokamaks to ITER. European Physical Journal H, 46(1):??, December 2021. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.

springer.com/article/10.1140/epjh/s13129-021-00009-6.

Paul:2013:YEP

[Pau13] Ewald Paul. 50 years of experimental particle physics in Bonn. A personal recollection. European Physical Journal H, 38(4):471–506, September 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2013-30028-7.

Pelogia:2017:AJI

[PB17] Karla Pelogia and Carlos Alexandre Brasil. Analysis of the Jun Ishiwara’s “The universal meaning of the quantum of action”. European Physical Journal H, 42(4–5):507–521, December 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/e2017-80034-x. See [Ish17].

Peebles:2014:DHB

[Pee14] Phillip James Edwin Peebles. Discovery of the hot Big Bang: What happened in 1948. European Physical Journal H, 39(2):205–223, April 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2014-50002-y.

Peebles:2017:RDN

[Pee17] Phillip James Edwin Peebles. Robert Dicke and the naissance of experimental gravity physics, 1957–1967. European Physical Journal H, 42(2):177–259, June 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

content/pdf/10.1140/epjh/e2016-70034-0.pdf.

Pellegrini:2012:HXR

[Pel12] C. Pellegrini. The history of X-ray free-electron lasers. European Physical Journal H, 37(5):659–708, October 2012. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2012-20064-5.

Penny:2013:SED

[Pen13] Alan John Penny. The SETI episode in the 1967 discovery of pulsars. European Physical Journal H, 38(4):535–547, September 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2012-30052-6.

Perryman:2012:HA

[Per12] Michael Perryman. The history of astrometry. European Physical Journal H, 37(5):745–792, October 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2012-30039-4.

Perkins:2013:ENE

[Per13] D. H. Perkins. An early neutrino experiment: how we missed quark substructure in 1963. European Physical Journal H, 38(5):713–726, December 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2013-40024-3.

Perkins:2014:RHD

[Per14] Don H. Perkins. The remarkable history of the discovery of neutrino oscillations. European Physical Journal H, 39(5):505–515, December 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2014-50037-4.

Perkins:2016:EST

[Per16] Don H. Perkins. Early steps towards quarks and their interactions using neutrino beams in CERN bubble chamber experiments. European Physical Journal H, 41(2):157–164, June 2016. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467

(electronic). URL http://link.springer.com/content/pdf/10.
1140/epjh/e2016-70016-2.pdf.

Pfister:2014:LLL

[Pfi14] Herbert Pfister. Ludwig Lange on the Law of Inertia. European Physical Journal H, 39(2):245–250, April 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2013-40039-2.

Perez:2022:IEO

[PI22] Enric Pérez and Joana Ibáñez. Indistinguishable elements in the origins of quantum statistics. The case of Fermi–Dirac statistics. European Physical Journal H, 47(1):??, December 2022. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-021-00033-6.

Picasso:2012:FMD

[Pic12] Emilio Picasso. A few memories from the days at LEP. European Physical Journal H, 36(4):551–562, March 2012. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2011-20050-0.

Pietschmann:2011:EHC

[Pie11] Herbert Pietschmann. The early history of current algebra. European Physical Journal H, 36(1):75–84, July 2011. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2011-20013-0.

Pizzella:2016:BID

[Piz16] G. Pizzella. Birth and initial developments of experiments with resonant detectors searching for gravitational waves. European Physical Journal H, 41(4–5):267–302, November 2016. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/accesspage/article/
10.1140/epjh/e2016-70036-8.

Plass:2012:CPS

[Pla12] Günther Plass. The CERN proton synchrotron: 50 years of reliable operation and continued development. European Physical Journal H, 36(4):439–454, March 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2011-20059-0.

Peliti:2023:RFP

[PMG23] Luca Peliti and Paolo Muratore-Ginanneschi. R. Fürth’s 1933 paper “On certain relations between classical statistics and quantum mechanics” [“Über einige Beziehungen zwischen klassischer Statistik und Quantenmechanik”, Zeitschrift für Physik, 81 143– 162]. European Physical Journal H, 48(1):??, December 2023. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-023-00052-5.

Perez:2016:BET

[PP16] Enric Pérez and Blai Pié i Valls. Bohr and Ehrenfest: transformations and correspondences in the early 1920s. European Physical Journal H, 41(2):93–136, June 2016. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.

springer.com/article/10.1140/epjh/e2016-60028-1.

Pagano:2019:NLT

[PP19] Angelo Pagano and Emanuele V. Pagano. A note on Lorentz transformations and simultaneity in classical physics and special relativity. European Physical Journal H, 44(4–5):321–330, November 2019. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2019-90058-4.

Peruzzi:2018:TPQ

[PR18] Giulio Peruzzi and Alessio Rocci. Tales from the prehistory of quantum gravity. European Physical Journal H, 43(2):185–241, May 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2018-80018-6.

Radvanyi:2013:DBP

[Rad13] Pierre Radvanyi. The discussion between P. Curie and E. Rutherford (1900–1904). European Physical Journal H, 38(4):433–441, September 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2013-30019-8.

Rickles:2015:PWG

[RB15] Dean Rickles and Alexander Blum. Paul Weiss and the genesis of canonical quantization. European Physical Journal H, 40(4– 5):469–487, December 2015. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2015-60001-5.

Ray:2014:ARG

[RBM14] Saibal Ray, Sudhindra Nath Biswas, and Utpal Mukhopadhyay. Astronomer R. G. Chandra: In the light of his Anglo–American connection. European Physical Journal H, 39(3):369–387, September 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2014-40019-6.

Rozelot:2011:HSO

[RD11] J.-P. Rozelot and C. Damiani. History of solar oblateness measurements and interpretation. European Physical Journal H, 36 (3):407–436, November 2011. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2011-20017-4.

Rozelot:2012:RWT

[RD12] J. P. Rozelot and C. Damiani. Rights and wrongs of the temporal solar radius variability. European Physical Journal H, 37 (5):709–743, October 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2012-20030-4.

Rebut:2018:JET

[Reb18] Paul-Henri Rebut. The Joint European Torus (JET). European Physical Journal H, 43(4–5):459–497, December 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2017-70068-y.

Reed:2024:RFP

[Ree24] B. Cameron Reed. Revisiting the Frisch–Peierls memorandum. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-024-00070-x.

Richter:2012:ECC

[Ric12] Burton Richter. Electron colliders at CERN. European Physical Journal H, 36(4):543–549, March 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2011-20056-0.

Rickles:2018:GWN

[Ric18] Dean Rickles. Geon Wheeler: from nuclear to spacetime physicist. European Physical Journal H, 43(3):243–265, August 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2018-80053-x.

Ray:2023:IGU

[RMD23] Saibal Ray, Utpal Mukhopadhyay, and Samir Dhurde. IUCAA: genesis of a unique research centre. European Physical Journal H, 48(1):??, December 2023. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.springer.

com/article/10.1140/epjh/s13129-023-00049-0.

Ray:2021:RSF

[RMS21] Saibal Ray, Utpal Mukhopadhyay, and Rajinder Singh. N. R. Sen: Father of Indian applied mathematics. European Physical Journal H, 46(1):??, December 2021. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.springer.

com/article/10.1140/epjh/s13129-021-00003-y.

Robinson:2019:GGR

[Rob19] D. C. Robinson. Gravitation and general relativity at King’s College London. European Physical Journal H, 44(3):181– 270, September 2019. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.

com/article/10.1140/epjh/e2019-100020-1;   http://link.
springer.com/content/pdf/10.1140/epjh/e2019-100020-1.pdf.

Rocci:2023:GEL

[Roc23] Alessio Rocci. On the gravitational effects of light. European Physical Journal H, 48(1):??, December 2023. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.

springer.com/article/10.1140/epjh/s13129-023-00055-2.

Rosenfeld:2017:QWF

[Ros17] Léon Rosenfeld. On the quantization of wave fields. European Physical Journal H, 42(1):63–94, April 2017. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/accesspage/article/

10.1140/epjh/e2016-70041-3. Originally published in German Zur Quantelung der Wellenfelder in Annalen der Physik 397, 113 (1930). Submitted for publication on March 18, 1930.

Rubbia:2012:PAC

[Rub12] Carlo Rubbia. Proton-antiproton colliders. European Physical Journal H, 36(4):523–542, March 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.

com/article/10.1140/epjh/e2011-20065-y.

Rocci:2024:QTG

[RV24] Alessio Rocci and Thomas Van Riet. The quantum theory of gravitation, effective field theories, and strings: yesterday and today. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-024-00069-4.

Salisbury:2022:HOH

[Sal22] Donald Salisbury. A history of observables and Hamilton–Jacobi approaches to general relativity. European Physical Journal H, 47(1):??, December 2022. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.springer.

com/article/10.1140/epjh/s13129-022-00039-8.

Sauer:2019:EWS

[Sau19] Tilman Sauer. Einstein’s working sheets and his search for a unified field theory. European Physical Journal H, 44(4–5):371–379, November 2019. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2019-100019-6.

Schmidt-Bocking:2016:SGE

[SBSL^{+}16] Horst Schmidt-Böcking, Lothar Schmidt, Hans Jürgen Lüdde, Wolfgang Trageser, Alan Templeton, and Tilman Sauer. The Stern– Gerlach experiment revisited. European Physical Journal H, 41(4– 5):327–364, November 2016. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

accesspage/article/10.1140/epjh/e2016-70053-2.

Schroer:2010:JSC

[Sch10] B. Schroer. Jorge A. Swieca’s contributions to quantum field theory in the 60s and 70s and their relevance in present research. European Physical Journal H, 35(1):53–88, July 2010. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2010-00004-1.

Schroer:2011:PJL

[Sch11] B. Schroer. Pascual Jordan’s legacy and the ongoing research in quantum field theory. European Physical Journal H, 35(4):377– 434, April 2011. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2011-10015-8.

Schopper:2012:CAE

[Sch12] Herwig Schopper. CERN’s accelerators, experiments and international integration 1959–2009. European Physical Journal H, 36(4): 437, March 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/content/
pdf/10.1140/epjh/e2012-20067-5.pdf.

Schroer:2013:EJC

[Sch13] Bert Schroer. The Einstein–Jordan conundrum and its relation to ongoing foundational research in local quantum physics. European Physical Journal H, 38(2):137–173, March 2013. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2012-30059-x.

Schrimpf:2014:FMD

[Sch14a] Andreas Schrimpf. The first measurement of the deflection of the vertical in longitude. European Physical Journal H, 39(3):389–402, September 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2014-40055-2.

Schrimpf:2014:ICC

[Sch14b] Andreas Schrimpf. An international campaign of the 19th century to determine the solar parallax. European Physical Journal H, 39(2):225–244, April 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2013-40036-2.

Schweber:2015:HQR

[Sch15] Silvan S. Schweber. Hacking the quantum revolution: 1925– 1975. European Physical Journal H, 40(1):53–149, March 2015. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2014-50060-9.

Scholz:2019:CAB

[Sch19] Erhard Scholz. E. Cartan’s attempt at bridge-building between Einstein and the Cosserats — or how translational curvature became to be known as torsion. European Physical Journal H, 44(1):47–75, February 2019. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2018-90059-x. See erratum [Sch20].

Scholz:2020:ECA

[Sch20] Erhard Scholz. Erratum to: E. Cartan’s attempt at bridgebuilding between Einstein and the Cosserats — or how translational curvature became to be known as torsion. European Physical Journal H, 45(4–5):345–374, December 2020. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https:

//link.springer.com/article/10.1140/epjh/e2020-0001s-y.

See [Sch19].

Sonin:2022:HZM

[SCS22] Anatoly S. Sonin, Natalia A. Churochkina, and Andrei A. Sonin. Hans Zocher and mineral liquid crystals. European Physical Journal H, 47(1):??, December 2022. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.springer.

com/article/10.1140/epjh/s13129-021-00035-4.

Sagdeev:2018:IRS

[SD18] Roald Z. Sagdeev and Patrick H. Diamond. An interview with Roald Sagdeev: his story of plasma physics in Russia, 1956– 1988. European Physical Journal H, 43(4–5):355–396, December 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2018-90042-3.

Smith:2020:MSC

[SGD20] William Smith, Martyn Guest, and Paul Durham. Molecular simulation and the collaborative computational projects. European Physical Journal H, 45(4–5):259–343, December 2020. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/e2020-10034-9.

Shim:2024:CTB

[Shi24] Jae Wan Shim. A commented translation of Boltzmann’s work, “Ueber die sogenannte H-Curve”. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.springer.

com/article/10.1140/epjh/s13129-024-00085-4.

Shlosman:2012:SMS

[Shl12] S. Shlosman. From the seminar on Mathematical Statistical Physics in Moscow State University, 1962–1994. Constructive criteria. European Physical Journal H, 37(4):595–603, September 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2012-10050-x.

Sinai:2012:SMS

[Sin12] Ya. G. Sinai. From the seminar on Mathematical Statistical Physics in Moscow State University, 1962–1994. How everything started. European Physical Journal H, 37(4):567–569, September 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2012-10055-6.

Smorenburg:2014:CFE

[SKL14] P. W. Smorenburg, L. P. J. Kamp, and O. J. Luiten. Classical formulations of the electromagnetic self-force of extended charged bodies. European Physical Journal H, 39(3):283–302, September 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2014-50015-2.

Stella:2011:GGD

[SM11] Bruno R. Stella and Hans-Jürgen Meyer. Υ (9.46 GeV) and the gluon discovery (a critical recollection of PLUTO results). European Physical Journal H, 36(2):203–243, September 2011. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2011-10029-3.

Smith:2013:WDM

[Smi13] Anders Smith. Who discovered the magnetocaloric effect? European Physical Journal H, 38(4):507–517, September 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2013-40001-9.

Smith:2023:CQQ

[Smi23] Chris Llewellyn Smith. From concrete quarks to QCD: a personal perspective. European Physical Journal H, 48(1):??, December 2023. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL https://link.springer.com/article/
10.1140/epjh/s13129-023-00061-4. See correction [Smi24].

Smith:2024:CCQ

[Smi24] Chris Llewellyn Smith. Correction to: From concrete quarks to QCD: a personal perspective. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.springer.

com/article/10.1140/epjh/s13129-024-00074-7. See [Smi23].

Soding:2010:DG

[Söd10] P. Söding. On the discovery of the gluon. European Physical Journal H, 35(1):3–28, July 2010. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2010-00002-5.

Sommerfeld:2014:FSH

[Som14a] A. Sommerfeld. The fine structure of hydrogen and hydrogenlike lines. European Physical Journal H, 39(2):179–204, April 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2013-40054-0.

Sommerfeld:2014:TBS

[Som14b] A. Sommerfeld. On the theory of the Balmer series. European Physical Journal H, 39(2):157–177, April 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2013-40053-8.

Spearman:2010:JM

[Spe10] T. D. Spearman. James MacCullagh 1809–1847. European Physical Journal H, 35(2):113–122, November 2010. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2010-00012-2.

Spiering:2012:CRG

[Spi12a] Christian Spiering. Cosmic rays, gamma rays and neutrinos: a survey of 100 years of research. European Physical Journal H, 37 (3):319–321, August 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

content/pdf/10.1140/epjh/e2012-30035-0.pdf.

Spiering:2012:THE

[Spi12b] Christian Spiering. Towards high-energy neutrino astronomy. European Physical Journal H, 37(3):515–565, August 2012. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2012-30014-2.

Salisbury:2017:LRG

[SS17] Donald Salisbury and Kurt Sundermeyer. Léon Rosenfeld’s general theory of constrained Hamiltonian dynamics. European Physical Journal H, 42(1):23–61, April 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.

com/content/pdf/10.1140/epjh/e2016-70042-7.pdf.

Steinhauser:2016:HSF

[Ste16] Georg Steinhauser. Hahn and Strassmann’s first credible, yet erroneous approximation to the discovery of nuclear fission. European Physical Journal H, 41(3):265–266, September 2016. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/accesspage/article/
10.1140/epjh/e2016-70043-y.

Straumann:2011:FSC

[Str11] Norbert Straumann. On the first Solvay Congress in 1911. European Physical Journal H, 36(3):379–399, November 2011. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2011-20043-9.

Schlatter:2012:SHL

[SZ12] W.-D. Schlatter and P. M. Zerwas. Searching for Higgs: from LEP towards LHC. European Physical Journal H, 36(4):579–600, March 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2012-20069-1.

Taltavull:2020:RLF

[Tal20] Marta Jordi Taltavull. Rudolf Ladenburg and the first quantum interpretation of optical dispersion. European Physical Journal H, 45(2–3):123–173, September 2020. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.

springer.com/article/10.1140/epjh/e2020-10027-6.

Tannenbaum:2016:WW

[Tan16] M. J. Tannenbaum. Waiting for the W. European Physical Journal H, 41(4–5):303–325, November 2016. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.

springer.com/accesspage/article/10.1140/epjh/e2016-70048-
7.

Tannenbaum:2018:HHC

[Tan18] M. J. Tannenbaum. How hadron collider experiments contributed to the development of QCD: from hard-scattering to the perfect liquid. European Physical Journal H, 43(2):119–183, May 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2017-80056-0.

Tipler:2013:PKA

[Tip13] Frank J. Tipler. Perry, Kelvin, and the age of the sun. European Physical Journal H, 38(3):405–409, April 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2012-30021-5.

Trimble:2010:OAC

[Tri10] Virginia Trimble. The origins and abundances of the chemical elements before 1957: from Prout’s hypothesis to Pasadena. European Physical Journal H, 35(1):89–109, July 2010. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2010-00006-9.

Trimble:2017:WW

[Tri17] Virginia Trimble. Wired by Weber. European Physical Journal H, 42(2):261–291, June 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

content/pdf/10.1140/epjh/e2016-70060-5.pdf.

Trautman:2019:MME

[TS19] Andrzej Trautman and Donald Salisbury. Memories of my early career in relativity physics. European Physical Journal H, 44(4–5):391–413, November 2019. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http:/

/link.springer.com/article/10.1140/epjh/e2019-100044-5;
 https://link.springer.com/content/pdf/10.1140/epjh/e2019-
100044-5.pdf.

Tsesmelis:2012:JSM

[Tse12] Emmanuel Tsesmelis. Jack Steinberger: Memories of the PS and of LEP. European Physical Journal H, 36(4):455–467, March 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2012-20066-1.

Taylor:2019:LPU

[TWZ19] Thomas Taylor, Horst Wenninger, and Antonino Zichichi. LAA: a project using dedicated funding to develop technology for highenergy physics experiments. European Physical Journal H, 44(4– 5):307–319, November 2019. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2019-100010-9.

vonNeumann:2010:PET

[vN10] J. von Neumann. Proof of the ergodic theorem and the Htheorem in quantum mechanics. European Physical Journal H, 35 (2):201–237, November 2010. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.com/

article/10.1140/epjh/e2010-00008-5.

Villone:2017:HHG

[VR17] Barbara Villone and Cornelius Rampf. Hermann Hankel’s “On the general theory of motion of fluids”. European Physical Journal H, 42(4–5):557–609, December 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic).

vanStrien:2021:WPE

[vS21] Marij van Strien. Was physics ever deterministic? The historical basis of determinism and the image of classical physics. European Physical Journal H, 46(1):??, December 2021. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-021-00012-x.

Vincent:2022:EOC

[VT22] Dwight E. Vincent and David R. Topper. Einstein’s Oxford cosmology blackboards: open portals to 1931. European Physical Journal H, 47(1):??, December 2022. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https://link.springer.

com/article/10.1140/epjh/s13129-022-00046-9.

Wagner:2018:HRI

[Wag18] F. Wagner. The history of research into improved confinement regimes. European Physical Journal H, 43(4–5):523–549, December 2018. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/
10.1140/epjh/e2016-70064-9;   http://link.springer.com/
content/pdf/10.1140/epjh%2Fe2016-70064-9.pdf.

Weidenmuller:2015:NPH

[Wei15] Hans A. Weidenmüller. Nuclear physics in Heidelberg in the years 1950 to 1980. Personal recollections. European Physical Journal H, 40(3):279–299, September 2015. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.springer.

com/article/10.1140/epjh/e2015-60019-4.

Weinberg:2021:DEF

[Wei21] Steven Weinberg. On the development of effective field theory. European Physical Journal H, 46(1):??, December 2021. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-021-00004-x.

Wu:2021:ETU

[Wu21] Biao Wu. Everett’s theory of the universal wave function. European Physical Journal H, 46(1):??, December 2021. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL

https://link.springer.com/article/10.1140/epjh/s13129-
021-00001-0.

Walter:2012:EHC

[WW12] M. Walter and A. W. Wolfendale. Early history of cosmic particle physics. European Physical Journal H, 37(3):323–358, August 2012. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2012-30020-1.

Zaghloul:2011:IFP

[Zag11] M. R. Zaghloul. Inconsistency in Fermi’s probability of the quantum states. European Physical Journal H, 36(3):401–406, November 2011. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2011-20018-8. See erratum [Zag13].

Zaghloul:2013:EIF

[Zag13] Mofreh R. Zaghloul. Erratum to: Inconsistency in Fermi’s probability of the quantum states. European Physical Journal H, 38(2):279, March 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467

(electronic). URL http://link.springer.com/content/pdf/10.
1140/epjh/e2013-40005-1.pdf. See [Zag11].

Zeh:2011:FIQ

[Zeh11] H. D. Zeh. Feynman’s interpretation of quantum theory. European Physical Journal H, 36(1):63–74, July 2011. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2011-10035-2.

Zinth:2011:LJL

[ZLK11] Wolfgang Zinth, Alfred Laubereau, and Wolfgang Kaiser. The long journey to the laser and its rapid development after 1960. European Physical Journal H, 36(2):153–181, September 2011. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2011-20016-0.